Hi im really stuck on this question:
m+10/6m=7/8 they are supposed to be fractions. Im really not sure how to do this. I would really appreciate your help
-Thanks Drea

Answer by stanbon(48568)

(m+10)/6m = 7/8
---
Cross-multiply to get:
8(m+10) = 7(6m)
8m + 80 = 42m
34m = 80
m = 40/17
-------------------
Cheers,
Stan H.